OpenStudy (anonymous):

Solve for x. 5(x + 1) = 4(x + 8)

OpenStudy (igreen):

Simplify: \(5(x + 1) = 4(x + 8)\) \(5x + 5 = 4x + 32\) \(5x = 4x + 27\) \(x = 27\)

OpenStudy (igreen):

You can give medals by clicking 'Best Response'. @i-need-help54